A small sphere of mass m is launched horizontally over a body of water from a height h above the water and with a launch speed v0. Determine expressions for the following in terms of m, v0, h, and g. Air resistance is negligibly small.

(a) W is the amount work done by the force of gravity on the projectile during its flight.

W =


(b) ΔKE is the change in kinetic energy ΔKE of the projectile from the time it was fired until it hits the water.

ΔKE =


(c) KEf is the final kinetic energy KEf of the projectile as it hits the water.

KEf =


(d) Are any of the answers changed if the initial angle is changed?
Answer #1

a)

W=9.81h

b)

t^{2}=\frac{2h}{g}

v_{y}=-gt=-\sqrt{2gh}

v_{x}=v_{0}

Final velocity (magnitude)

v_{f}=\sqrt{2gh+v_{0}^{2}}

Change in Kinetic energy:

\Delta KE=\frac{1}{2}m(v_{0}^{2}-2gh-v_{0}^{2})

\Delta KE=-\frac{1}{2}m(2gh)

-----------------------------------------------

Final Kinetic energy:

KEf=\frac{1}{2}m(2gh+v_{0}^{2})